# Backdoor Attack with Sparse and Invisible Trigger

### v2

@@ I Introduction
However, the convenience of using third-party training resources may also lead to new security threats. One of the most emerging yet severe threats is called the backdoor attack [4, 5, 6, 7, 8, 9], where the adversaries intend to implant a hidden backdoor to victim models during the training process. The backdoored models behave normally on predicting benign samples whereas their predictions will be maliciously changed to the adversary-specified target class whenever the backdoor is activated by the adversary-specified trigger pattern. Backdoor attacks severely reduced the model’s reliability and attracted tremendous attention from various domains. A recent industry report [10] found that companies are concerned about backdoor attacks and ranked them as the fourth among the most popular security threats. Besides, government agencies also placed a high priority on backdoor research. For instance, U.S. intelligence community [11] launched a new funding program to defend against backdoor attacks and some other threats.
@@
The design of trigger patterns is one of the most important factors in backdoor attacks. Currently, there are many different types of trigger patterns [12, 13, 9]. Arguably, patch-based triggers [4, 14, 15] and additive triggers [16, 17, 18] are the most classical and widely adopted ones among all trigger patterns. Specifically, patch-based triggers are extremely sparse since they only modify a small number of pixels. Accordingly, they can be implemented as stickers, making the attacks more feasible in the physical world. However, they are visible to human eyes and can not evade inspection. On the other hand, additive triggers are usually invisible but modify almost all pixels, restricting the feasibility in real scenarios. An intriguing question arises: Is it possible to design a trigger pattern for effective backdoor attacks that is both sparse and invisible?
@@
The answer to the aforementioned question is positive. In general, the most straightforward method is to introduce a (random) sparse mask to the patterns of existing invisible backdoor attacks during their generation process. However, as we will show in the experiments, this method is not effective in many cases, especially when the image size is relatively large. Its failure is mostly because the position of trigger areas is also critical for their success, especially when the trigger size and perturbation strength are limited. Based on these understandings, in this paper, we propose to design sparse and invisible backdoor attacks by optimizing trigger areas and patterns simultaneously. Specifically, we formulate it as a bi-level optimization problem with sparsity and invisibility constraints. The upper-level problem is to minimize the loss on poisoned samples via optimizing the trigger, while the lower-level one is to minimize the loss on all training samples via optimizing the model weights. In particular, this optimization problem is difficult to solve directly due to the high non-convexity and poor convergence of gradient-based methods. To alleviate these problems, we exploit a pre-trained surrogate model to reduce the complexity of lower-level optimization and derive an alternate projected method to satisfy the $L_{\infty}$ and $L_{0}$ constraints. We conduct extensive experiments on benchmark datasets to demonstrate the effectiveness of the proposed method and its feasibility under different settings. Besides, as we will show in the experiments, the generated sparse trigger pattern contains semantic information about the target class. It indicates our attack may serve as the potential path toward explaining DNNs, which is an unexpected gift.
@@ II-B Backdoor Defense
Currently, there are some methods (dubbed ‘backdoor defenses’) to reduce the backdoor threats. In general, existing defenses can be divided into four main categories, including (1) the detection of poisoned training samples, (2) poison suppression, (3) backdoor removal, and (4) the detection of poisoned testing samples. They correspond to the four important stages of DNNs, including data collection, model training, model deployment, and model inference.
Specifically, the first type of defense intends to filter out poisoned training samples [24, 25, 26]. They are either based on the representation differences between poisoned and benign samples at intermediate layers of the model or directly attempt to reverse the trigger pattern; Poison suppression [27, 28, 29] depresses the effectiveness of poisoned samples during the training process to prevent the creation of hidden backdoors; Backdoor removal aims to remove the hidden backdoors in give pre-trained models [30, 31, 32]; The last type of defense [33, 34, 35] targets the detection of poisoned testing samples, serving as a firewall of given models. These defenses are probably the most practical ones, due to their efficiency and low requirements of defender capacities.
@@ III-A Preliminaries
Threat Model. In this paper, we focus on the poison-only backdoor attacks, where the adversaries can only manipulate training samples without having any information about the victim model. We study the image classification task where the model outputs a $C$-class probability vector: $f_{\bm{\theta}}:\mathbb{R}^{d}\rightarrow[0,1]^{C}$. Given a training set $\mathcal{D}=\left\{{\bm{x}}_{i},y_{i}\right\}_{i=1}^{N}$, $\bm{x}\in\mathbb{R}^{d}$, $y\in\left[C\right]=[1,2,\cdots,C]$, the parameters of classifier are optimized by the empirical risk minimization:
$\min\sum_{({\bm{x}},y)\in\mathcal{D}}\mathcal{L}(f_{\bm{\theta}}({\bm{x}}),y),$
where $\mathcal{L}$ represents the loss function ($e.g.$, the cross entropy loss). The adversary is allowed to modify a small subset $\mathcal{D}_{s}$ of the original training set $\mathcal{D}$ ($i.e.$, $\mathcal{D}_{s}\subset\mathcal{D}$ and $|\mathcal{D}_{s}|\ll|\mathcal{D}|$) by attaching the trigger $\bm{t}$ and relabelling them as the target class $y_{T}$. The victim model $f_{\bm{\theta}}$ is trained on the modified dataset $\mathcal{D}^{\prime}$, which is composed of a benign dataset $\mathcal{D}_{c}=\mathcal{D}\setminus\mathcal{D}_{s}$ and a poisoned dataset $\mathcal{D}_{p}=\left\{({\bm{x}}+{\bm{t}},y_{T})|({\bm{x}},y)\in\mathcal{D}_{s}\right\}$. In particular, the ratio $|\mathcal{D}_{p}|/|\mathcal{D}|$ is called as the poisoning rate.
Adversary Goals. In general, the adversaries have three main goals, including the utility, effectiveness, and stealthiness. Specifically, utility requires that the attacked model $f_{\bm{\theta}}$ achieves high accuracy on benign test samples. Otherwise, the model would not be adopted and no backdoor could be planted; The effectiveness desires that the attacked model can achieve high attack success rates when benign samples are attached with the trigger pattern; The stealthiness requires that the dataset modification should be unnoticeable to victim dataset users. For example, the trigger patterns should be invisible and the poisoning rate should be small.
@@ III-B Sparse and Invisible Backdoor Attack (SIBA)
To solve Problem 4, we have the following Lemma.
Lemma 1.
[…]
@@
IV Experiments
IV-A Experimental Setup
Attack Settings. For all attacks, we set the poisoning rate as $1\%$ and choose the class ‘0’ as the target class. Specifically, for the settings of BadNets, the trigger is a $3\times 3$ checkerboard on CIFAR-10 and a $20\times 20$ all-white patch on VGGFace2; The trigger is a Hello-Kitty image on CIFAR-10 and a random noise image on VGGFace2 for Blended. The transparency parameter is set as $0.2$; We exploit a pre-trained model to generate the targeted universal adversarial perturbation on the benign model as the trigger for TUAP. The $L_{\infty}$ constraint is set as $8/255$ on both datasets; For WaNet, the size of the control field is $4\times 4$ and the strength parameter of the backward warping field is set to $0.5$; We adopt the default settings used in its original paper [18] for ISSBA; UBW-P is an untargeted backdoor with the BadNets-type trigger pattern; For our SIBA, we set the step size $\alpha=0.2$ and $L_{\infty}$ constraint $\epsilon=8/255$ on both datasets. $L_{0}$ constraint $k$ is set to $100$ on CIFAR-10 dataset and $1,600$ on VGGface2 dataset. We set the number of training iterations $T=200$ and the update step $K=5$ on both datasets. We use the whole training set to optimize the SIBA trigger. The example of poisoned samples is shown in Figure 2.
@@ IV-D The Resistance to Potential Backdoor Defenses
The Resistance to Neural Cleanse. As a representative detection of poisoned training samples, neural cleanse (NC) [25] first attempts to reconstruct the possible trigger of the candidate model and collects the $L_{1}$ values of the optimized trigger associated with each target label. Then, NC calculates the median absolute deviation of the group and the anomaly index of each label. If the anomaly index is larger than the threshold, the model is regarded as backdoored and the defender can filter out poisoned training samples. In our experiments, the threshold is 2 as suggested in its original paper. We use the Adam optimizer in which the learning rate is $0.1$. The coefficient of the regularization term is $0.001$ and the number of training epochs is 50. As shown in Figure 7, the NC is ineffective to detect our backdoored model since the anomaly index of the proposed method is always lower than the threshold value. This failure is mostly because our SIBA only needs to manipulate a small number of pixels such that the optimization of NC cannot catch our trigger location.
@@
Figure 7: The resistance of our SIBA to neural cleanse.
TABLE III: The resistance to anti-backdoor learning.
@@
The Resistance to Fine-pruning. As a representative backdoor removal method, fine-pruning (FP) [30] first tests the candidate model with a small clean validation set and records the average activation of each neuron. Then, FP prunes the channels with increasing order until the clean accuracy drops below some threshold. In our experiments, the validation set is obtained by randomly choosing $20\%$ samples from the clean training dataset and the total channel number is 512. The curves of BA and ASR with respect to the number of pruned channels are shown in Figure 9. We could observe that the ASR of the proposed backdoor attack preserves on CIFAR-10 even if a large portion of channels are pruned. As for the VGGFace2 dataset, the ASR is reduced below $80\%$ when the number of pruned channels is larger than $400$. However, the BA is significantly decreased as its sacrifice. These results verify the resistance of our SIBA to FP.
@@
Figure 9: The resistance of our SIBA to FP.
(a) CIFAR-10 (ResNet)
(b) CIFAR-10 (VGG)
(c) VGGFace2 (ResNet)
(d) VGGFace2 (VGG)
Figure 10: The resistance of our SIBA to STRIP.
@@
The Resistance to STRIP. As a representative black-box detection of poisoned samples with predicted logits, STRIP [34] perturbs a given test image by superimposing various images and then inspects the entropy of the model prediction. The suspicious samples having low entropy are regarded as poisoned samples. We evaluate the resistance of our SIBA to STRIP by visualizing the entropy distributions of samples. As shown in Figure 10, the entropy distributions of poisoned samples are mixed with those of benign samples. Accordingly, our SIBA can evade the detection of STRIP.
TABLE IV: The resistant of our SIBA to Scale-Up.
@@
The Resistance to Scale-Up. As a representative black-box detection of poisoned samples with predicted labels, Scale-Up [35] discovered the phenomenon that the poisoned samples had the scaled prediction consistency when the pixel values were amplified and proposed to distinguish the poisoned samples by counting the predictions of scaled images. In our experiments, we use a scaling set with size 5 and set the threshold as $0.8$. We report true positive rate (TPR), false positive rate (FPR), area under the receiver operating characteristic (AUROC), and ASR in Table IV. As shown in the table, although Scale-Up can decrease the effectiveness of SIBA to some extent, the detection performance is far from satisfactory since the average ASR is still $>70\%$. In other words, our SIBA is resistant to the Scale-Up to a large extent.
@@ IV-E1 A Closer Look to the Effectiveness of our SIBA and its Connection to DNN Interpretability
To understand the effectiveness of our SIBA, we also visualize the (normalized) trigger patterns when different target labels are adopted. As shown in Figure 12, the SIBA trigger is always located in the main body of the object from the target class (such as the airplane’s fuselage and wings in Figure 12). In other words, the generated sparse trigger pattern is closely related to the concept of the target label. This phenomenon also (partially) explains the prediction and learning mechanisms of DNNs. Specifically, from the prediction perspective, it indicates that the misclassification of DNNs towards poisoned samples from their ground-truth label to the target one might be attributed to perturbations in regions associated with the target class. From the prediction perspective, it implies that DNNs learn the trigger pattern as the representative of samples from the target class since we only re-assign the label of selected samples as the target label when generating the trigger pattern. It suggests that our attack may be a viable path toward understanding the learning and prediction principles of DNNs. Our attack builds a bridge between two seemingly unrelated fields ($i.e.$, backdoor attacks and DNN interpretability), which is an unexpected gift.

### v3

@@ I Introduction
However, the convenience of using third-party training resources may also lead to new security threats. One of the most emerging yet severe threats is called the backdoor attack [4, 5, 6, 7, 8, 9], where the adversaries intend to implant a hidden backdoor to victim models during the training process. The backdoored models behave normally on predicting benign samples whereas their predictions will be maliciously changed to the adversary-specified target class whenever the backdoor is activated by the adversary-specified trigger pattern. Backdoor attacks severely reduce the model’s reliability and attract tremendous attention from various domains. A recent industry report [10] demonstrated that companies are concerned about backdoor attacks and rank them as the fourth among the most popular security threats. Besides, government agencies also placed a high priority on backdoor research. For instance, U.S. intelligence community [11] launched a new funding program to defend against backdoor attacks and some other threats.
@@
The design of trigger patterns is one of the most important factors in backdoor attacks. Currently, there are many different types of trigger patterns [12, 13, 9]. Arguably, patch-based triggers [4, 14, 15] and additive triggers [16, 17, 18] are the most classical and widely adopted ones among all trigger patterns [^1] The trigger patterns vary depending on the specific tasks. In this work, we focus on image classification tasks. . Specifically, patch-based triggers are extremely sparse since they only modify a small number of pixels. Accordingly, they can be implemented as stickers, making the attacks more feasible in the physical world. However, they are visible to human eyes and can not evade inspection. On the other hand, additive triggers are usually invisible but modify almost all pixels, restricting the feasibility in real scenarios. Overall, it is crucial for a backdoor trigger to maintain both sparseness and imperceptibility since the sparseness benefits the practical implementation while imperceptibility helps to evade immediate detection and post hoc forensic analysis. However, previous works [4, 16, 17, 18, 14, 15] only satisfy at most one requirement. An intriguing question arises: Is it possible to design a trigger pattern for effective backdoor attacks that is both sparse and invisible?
@@
The answer to the aforementioned question is positive. In general, the most straightforward method is to introduce a (random) sparse mask to the patterns of existing invisible backdoor attacks during their generation process. However, as we will show in the experiments, this method is not effective in many cases, especially when the image size is relatively large. Its failure is mostly because the position of trigger areas is also critical for their success, especially when the trigger size and perturbation strength are limited. Based on these understandings, in this paper, we propose to design sparse and invisible backdoor attacks by optimizing trigger areas and patterns simultaneously. Specifically, we formulate it as a bi-level optimization problem with sparsity and invisibility constraints. The upper-level problem is to minimize the loss on poisoned samples via optimizing the trigger, while the lower-level one is to minimize the loss on all training samples via optimizing the model weights. In particular, this optimization problem is difficult to solve directly due to the high non-convexity and poor convergence of gradient-based methods. To alleviate these problems, we exploit a pre-trained surrogate model to reduce the complexity of lower-level optimization and derive an alternate projected method to satisfy the $L_{\infty}$ and $L_{0}$ constraints. We conduct extensive experiments on benchmark datasets to demonstrate the effectiveness of the proposed method and its feasibility under different settings. Besides, as we will show in the experiments, the generated sparse trigger pattern contains semantic information about the target class. It indicates our attack may serve as the potential path toward explaining DNNs, which is an unexpected finding.
Our main contributions can be summarized as follows.
@@ II-A Backdoor Attack
Recently, there were also a few works exploiting and designing backdoor attacks for positive purposes [24, 25, 26, 27], which are out of the scope of this paper.
@@ II-B Backdoor Defense
Currently, there are some methods (dubbed ‘backdoor defenses’) to reduce the backdoor threats. In general, existing defenses can be divided into five main categories, including (1) the detection of poisoned training samples, (2) poison suppression, (3) backdoor removal, (4) the detection of poisoned testing samples, and (5) the detection of attacked models.
Specifically, the first type of defense intends to filter out poisoned training samples [28, 29, 30]. They are either based on the representation differences between poisoned and benign samples at intermediate layers of the model or directly attempt to reverse the trigger pattern; Poison suppression [31, 32, 33] depresses the effectiveness of poisoned samples during the training process to prevent the creation of hidden backdoors; Backdoor removal aims to remove the hidden backdoors in give pre-trained models [34, 35, 36]; The forth type of defense [37, 38, 39] targets the detection of poisoned testing samples, and the last type of defense determines whether a given model is backdoored based on some model properties [29, 40, 41].
@@ III-A Preliminaries
Threat Model. In this paper, we study the image classification task where the model outputs a $C$-class probability vector: $f_{\bm{\theta}}:\mathbb{R}^{d}\rightarrow[0,1]^{C}$. Given a training set $\mathcal{D}=\left\{{\bm{x}}_{i},y_{i}\right\}_{i=1}^{N}$, $\bm{x}\in\mathbb{R}^{d}$, $y\in\left[C\right]=\left\{1,2,\cdots,C\right\}$, the parameters of classifier are optimized by the empirical risk minimization:
$\min_{\bm{\theta}}\sum_{({\bm{x}},y)\in\mathcal{D}}\mathcal{L}(f_{\bm{\theta}}({\bm{x}}),y),$
where $\mathcal{L}$ represents the loss function ($e.g.$, the cross entropy loss). Following the analysis framework in [59], we demonstrate our threat model from four perspectives: adversary’s goal, knowledge, capability, and strategy, as follows.
Adversary’s Goal. In general, the adversaries have three main goals, including the utility, effectiveness, and stealthiness. Specifically, the utility requires that the attacked model $f_{\bm{\theta}}$ achieves high accuracy on benign test samples. Otherwise, the model would not be adopted and therefore no backdoor could be implanted; The effectiveness desires that the attacked model can achieve high attack success rates whenever trigger patterns appear; The stealthiness requires that the dataset modification should be unnoticeable to victim dataset users. For example, the trigger patterns should be invisible and sparse, while the poisoning rate should be small.
Adversary’s Knowledge. We assume the adversary has access to (a few) training data but neither the learning algorithm nor the objective function during the training. Our settings align with previous works on backdoor attacks [4, 16, 18] and resemble the limited-knowledge gray-box attacks proposed in [59].
Adversary’s Capability. The adversary is only allowed to modify a small subset $\mathcal{D}_{s}$ of the original training set $\mathcal{D}$ ($i.e.$, $\mathcal{D}_{s}\subset\mathcal{D}$ and $|\mathcal{D}_{s}|\ll|\mathcal{D}|$) by attaching the trigger $\bm{t}$ and relabelling them as the target class $y_{T}$. The victim model $f_{\bm{\theta}}$ is trained on the modified dataset $\mathcal{D}^{\prime}$, which is composed of a benign dataset $\mathcal{D}_{c}=\mathcal{D}\setminus\mathcal{D}_{s}$ and a poisoned dataset $\mathcal{D}_{p}=\left\{({\bm{x}}+{\bm{t}},y_{T})|({\bm{x}},y)\in\mathcal{D}_{s}\right\}$. In particular, the ratio $|\mathcal{D}_{p}|/|\mathcal{D}|$ is called as the poisoning rate.
Attack Strategy. We formulate the proposed attack as a bi-level optimization. Its details are in the next subsection.
@@ III-B Sparse and Invisible Backdoor Attack (SIBA)
To solve Problem 4, we have the following Lemma.
Lemma 1.
${\bm{t}}_{i+1,j}=\left\{\begin{aligned} &{\bm{v}}_{i,j}&\text{if}\ j\in C^{\prime}\\ &0&\text{if}\ j\notin C^{\prime}\end{aligned},\right.$ (5)
@@
IV Experiments
IV-A Experimental Setup
Attack Settings. For all attacks, we set the poisoning rate as $1\%$ and choose the class ‘0’ as the target class. Specifically, for the settings of BadNets, the trigger is a $3\times 3$ checkerboard on CIFAR-10 and a $20\times 20$ all-white patch on VGGFace2; The trigger is a Hello-Kitty image on CIFAR-10 and a random noise image on VGGFace2 for Blended. The transparency parameter is set as $0.2$; We exploit a pre-trained model to generate the targeted universal adversarial perturbation on the benign model as the trigger for TUAP. The $L_{\infty}$ constraint is set as $8/255$ on both datasets; For WaNet, the size of the control field is $4\times 4$ and the strength parameter of the backward warping field is set to $0.5$; We adopt the default settings used in its original paper [18] for ISSBA; UBW-P is an untargeted backdoor with the BadNets-type trigger pattern; For our SIBA, we set the step size $\alpha=0.2$ and $L_{\infty}$ constraint $\epsilon=8/255$ on both datasets. $L_{0}$ constraint $k$ is set to $100$ on CIFAR-10 dataset and $1,600$ on VGGface2 dataset. We set the number of training iterations $T=200$ and the update step $K=5$ on both datasets. We use the whole training set to optimize the SIBA trigger. We implement them based on BackdoorBox [66]. The example of poisoned samples is shown in Figure 2.
@@ IV-D The Resistance to Potential Backdoor Defenses
The Resistance to STRIP. As a representative black-box detection of poisoned training samples with predicted logits, STRIP [38] perturbs a given test image by superimposing various images and then inspects the entropy of the model prediction. The suspicious samples having low entropy are regarded as poisoned samples. We evaluate the resistance of our SIBA to STRIP by visualizing the entropy distributions of samples. As shown in Figure 8, the entropy distributions of poisoned samples are mixed with those of benign samples. Accordingly, our SIBA can evade the detection of STRIP.
@@
TABLE III: The resistance to anti-backdoor learning (ABL).
@@
The Resistance to Fine-pruning (FP). As a representative backdoor removal method, fine-pruning (FP) [34] first tests the candidate model with a small clean validation set and records the average activation of each neuron. Then, FP prunes the channels with increasing order until the clean accuracy drops below some threshold. In our experiments, the validation set is obtained by randomly choosing $20\%$ samples from the clean training dataset and the total channel number is 512. The curves of BA and ASR with respect to the number of pruned channels are shown in Figure 9. We could observe that the ASR of the proposed backdoor attack preserves on CIFAR-10 even if a large portion of channels are pruned. As for the VGGFace2 dataset, the ASR is reduced below $80\%$ when the number of pruned channels is larger than $400$. However, the BA is significantly decreased as its sacrifice. These results verify the resistance of our SIBA to FP.
Figure 8: The resistance of our SIBA to STRIP.
@@
Figure 9: The resistance of our SIBA to fine-pruning (FP).
@@
Figure 11: The resistance of our SIBA to neural cleanse (NC).
TABLE IV: The resistance of our SIBA to Scale-Up.
@@
The Resistance to Scale-Up. As a representative black-box detection of poisoned testing samples with predicted labels, Scale-Up [39] discovered the phenomenon that the poisoned samples had the scaled prediction consistency when the pixel values were amplified and proposed to distinguish the poisoned samples by counting the predictions of scaled images. In our experiments, we use a scaling set with size 5 and set the threshold as $0.8$. We report true positive rate (TPR), false positive rate (FPR), area under the receiver operating characteristic (AUROC), and ASR in Table IV. As shown in the table, although Scale-Up can decrease the effectiveness of SIBA to some extent, the detection performance is far from satisfactory since the average ASR is still $>70\%$. In other words, our SIBA is resistant to the Scale-Up to a large extent.
@@
The Resistance to Neural Cleanse (NC). As a representative model-level detection method, NC [29] first reverses possible triggers of the suspicious model and collects the $L_{1}$ values of the optimized trigger associated with each target label. Then, NC calculates the median absolute deviation of the group and the anomaly index of each label. If the anomaly index is larger than the threshold, the model is regarded as backdoored. In our experiments, the threshold is 2 as suggested in its original paper. We use the Adam optimizer in which the learning rate is $0.1$. The coefficient of the regularization term is $0.001$ and the number of training epochs is 50. As shown in Figure 11, the NC is ineffective to detect our backdoored model since the anomaly index of the proposed method is always lower than the threshold value. This failure is mostly because our SIBA only needs to manipulate a small number of pixels such that the optimization of NC cannot catch our trigger location.
The Resistance to Meta Neural Trojan Detection (MNTD). The MNTD [40] is another famous model-level detection method. It trains a meta-classifier to determine whether a target model is backdoored. In our experiments, we set the query number 10 and use Adam optimizer [70] with a 0.001 learning rate to train the meta classifier. Besides, we train 20 candidate models (10 SIBA-backdoored models and 10 clean models) and then calculate the trojan scores based on its meta classifier. The result (AUROC=$0.57$) indicates that SIBA can escape from MNTD. It is mostly because MNTD cannot capture sufficient trigger-related features due to our trigger sparseness. We will study it further in our future works.
@@ IV-E1 A Closer Look to the Effectiveness of our SIBA and its Connection to DNN Interpretability
To understand the effectiveness of our SIBA, we also visualize the (normalized) trigger patterns when different target labels are adopted. As shown in Figure 12, the SIBA trigger is always located in the main body of the object from the target class (such as the airplane’s fuselage and wings in Figure 12). In other words, the generated sparse trigger pattern is closely related to the concept of the target label. This phenomenon also (partially) explains the prediction and learning mechanisms of DNNs. Specifically, from the prediction perspective, it indicates that the misclassification of DNNs towards poisoned samples from their ground-truth label to the target one might be attributed to perturbations in regions associated with the target class. From the learning perspective, it implies that DNNs learn the trigger pattern as the representative of samples from the target class since we only re-assign the label of selected samples as the target label when generating the trigger pattern. It suggests that our attack may be a viable path toward understanding the learning and prediction principles of DNNs. Our results build an interesting connection between backdoor attacks and DNN interpretability that has not been shown before.